# Grasp as You Say: Language-guided Dexterous Grasp Generation

### v1

@@ Abstract
This paper explores a novel task “Dexterous Grasp as You Say” (DexGYS), enabling robots to perform dexterous grasping based on human commands expressed in natural language. However, the development of this field is hindered by the lack of datasets with natural human guidance; thus, we propose a language-guided dexterous grasp dataset, named DexGYSNet, offering high-quality dexterous grasp annotations along with flexible and fine-grained human language guidance. Our dataset construction is cost-efficient, with the carefully-design hand-object interaction retargeting strategy, and the LLM-assisted language guidance annotation system. Equipped with this dataset, we introduce the DexGYSGrasp framework for generating dexterous grasps based on human language instructions, with the capability of producing grasps that are intent-aligned, high quality and diversity. To achieve this capability, our framework decomposes the complex learning process into two manageable progressive objectives and introduce two components to realize them. The first component learns the grasp distribution focusing on intention alignment and generation diversity. And the second component refines the grasp quality while maintaining intention consistency. Extensive experiments are conducted on DexGYSNet and real world environment for validation.
@@ 1 Introduction
In this paper, we explore a novel task "Dexterous Grasp as You Say", term DexGYS, as shown in Figure 1. We can see that natural human guidance is provided in this task, and can be utilized to drive dexterous grasping generation, thereby facilitating more user-friendly human-robot interactions. However, the new task also brings in new challenges. First, the high costs of annotating dexterous pose and the corresponding language guidance, present a barrier for developing and scaling dexterous datasets. Second, the demands of generating dexterous grasps that ensure intention alignment, high quality and diversity, present considerable challenges to the model learning.
To address the first challenge, we propose a large-scale language-guided dexterous grasping dataset DexGYSNet. DexGYSNet is constructed in a cost-effective manner by exploiting human grasp behavior and the extensive capabilities of Large Language Models (LLM). Specially, we introduce the Hand-Object Interaction Retargeting (HOIR) strategy to transfer easily-obtained human hand-object interactions to robotic dexterous hand, to maintain contact consistency and high-quality grasp posture. Subsequently, we develop the LLM-assisted Language Guidance Annotation system to produce flexible and fine-grained language guidance for dexterous grasp data with the support of LLM. DexGYS dataset comprises 50,000 pairs of high-quality dexterous grasps and their corresponding language guidance, on 1,800 common household objects.
@@
With the support of the dataset, we now turn our way to overcome the second challenge. And we propose the DexGYSGrasp framework for dexterous grasp generation, which aligns with intentions, ensures high quality, and maintains diversity. At the beginning, we surprisingly find the difficulty of mastering all objectives simultaneously results from the commonly used penetration loss unidexgrasp ; xu2024dgtr , which used to avoid hand-object penetration. As shown in Figure 2, penetration loss substantially hinders the learning of grasp distribution, causing intention misalignment and reduced diversity. Conversely, despite the high diversity and aligned intention, the removal of penetration loss leads to unacceptable object penetration, making the grasp infeasible. Based on this finding, we design our DexGYSGrasp framework in a progressive strategy, decomposing the complex learning task into two sequential objectives managed by progressive components. Initially, the first components learns a grasp distribution, which focuses on intention consistency and diversity, optimizing effectively without the constraints of penetration loss. Subsequently, the second components refines the initial coarse grasps to high-quality ones with the same intentions and diversity. Our framework allows each component to focus on specific and manageable optimization objective, significantly enhancing the overall performance of the generated grasps.
@@ 2.1 Dexterous Grasp Generation
Dexterous hand endows robots with the capability to manipulate objects in a human-like manner. Previous methods have achieved impressive results in ensuring grasp stability by analytical approaches analytic_1 ; DFC ; q1 ; dexgraspnet ; graspd and deep learning methods lu2023ugg ; unigrasp ; xu2024dgtr ; unidexgrasp ; efficientgrasp ; generating_multi_finger . However, the full potential of dexterous hands for intentional and human-like grasping has not been completely exploited in these methods. Recently, some works have focused on functional dexterous grasping chen2023TaskDex ; zhu2023FunctionDex2 ; wei2023FunctionDex1 ; zhu2021toward , aiming to achieve human-like capabilities that extend beyond grasp stability alone, but are still lack of flexibility and generalization. In this work, we explore a novel task of Dexterous Grasp as You Say, which fully leverages the dexterity of robotic hands and enable robot to execute dexterous grasp based on human natural language.
@@ 2.2 Grasp Datasets
The development of large-scale datasets has significantly improve the advancement of data-driven grasp methods, including parallel grasp fang20201billion ; eppner2021acronym ; murali2021task2_1 ; tang2023task2_2 , human grasp hasson2019obman ; chao2021dexycb ; yang2022oakink ; jian2023affordpose ; li2024semgrasp ; wang2024s2p , and dexterous grasp approaches ddg ; gendexgrasp ; chen2023TaskDex ; dexgraspnet ; unidexgrasp . Despite these advancements, the high cost of data collection remains a significant challenge, particularly in the domain of dexterous hands. Previous datasets for dexterous grasping primarily rely on physical analysis approaches DFC ; miller2004graspit to mitigate this issue. However, these approaches often lack the specific semantic context or corresponding language guidance necessary for constructing our language-guided dexterous task. In this paper, we present DexGYSNet dataset, with a cost-effective construction, providing high-quality dexterous grasp annotation along with flexible and fine-grained human language guidance.
@@ 2.3 Language-guided Robot Grasp
Language-guided robot grasp is important in robotics. Previous works focusing on parallel grippers have made strides in achieving task-oriented graspingmurali2021task2_1 ; tang2023task2_2 ; tang2023graspgpt , language-guided graspingxu2023joint ; jin2024reasoning and manipulation guided by languagejang2022bc ; mees2022matters ; driess2023palme ; shridhar2021cliport . In contrast to parallel grippers, dexterous hand boast a higher number of DOF (e.g., 28 for the Shadow Handshadowhand ), enabling a broader dexterity. However, this high freedom also presents challenges for model learning. In this paper, we propose a DexGYSGrasp framework, capable of generating intention-aligned dexterous grasps with high-quality and diversity.
@@ 3.2 Hand-Object Interaction Retargeting
Our Hand-Object Interaction Retargeting (HOIR) aims to transfer human hand-object interaction to dexterous hand-object interaction as shown in Figure 3 . The source MANO romero2017MANO hand parameters are denoted as $\mathcal{G}^{m}\in{R}^{61}$. And the target dexterous hand parameters are denoted as $\mathcal{G}^{dex}$ = $(r,t,q)$, where $r\in\mathbf{SO(3)}$ represents the global rotation, $t\in\mathbb{R}^{3}$ is the translation in world coordinates, and $q\in\mathbb{R}^{J}$ is the joint angles for a $J$-DoF dexterous hand, for example $J=22$ for Shadow Handshadowhand
Three steps are within the HOIR: pose initialization, fingertip alignment, and interaction refinement. In the first step, the dexterous poses are initialized by copying parameters from similar structures of human poses to establish better initial values. In the second step, the dexterous poses are optimized in the parameter space to align the fingertip positions $p_{k}^{dex,ft}$ with those of the human $p_{k}^{mano,ft}$. This achieve retargeting consistency, and the optimization objective can be formulated as follows:
@@
To improve the physical interaction feasibility while maintaining the consistency, the dexterous hand poses are further optimized in the third step by hand-object interaction and physical constraints losses dexgraspnet . Two key points are designed to maintain the consistency: preserving the contact area of the optimized pose consistent with the output from the second step, and keeping the translation fixed during this phase. The optimization objective can be formulated as follows:
@@
Here, the object penetration loss $\mathcal{L}_{pen}$ penalizes the depth of hand-object penetration. The self-penetration loss $\mathcal{L}_{spen}$ penalizes the self-penetration The joint angle loss $\mathcal{L}_{spen}$ penalizes the out-of-limit joint angles. The contact map loss $\mathcal{L}_{cmap}$ ensures the contact map on the object remains consistent with the output from the second stage.
@@ 3.3 LLM-assisted Language Guidance Annotation
To annotate flexible and fine-grained language guidance for dexterous hand-object pairs at a low cost, we design a coarse-to-fine automated language guidance annotation system with the assistance of the LLM, inspired by cui2024anyskill ; li2024semgrasp , as shown in Figure 3. Specially, we initially generate brief guidance based on the object category and brief human intentions (e.g., "using a lotion pump"), which are collected by the human dataset yang2022oakink . Subsequently, we compile the contact information for each finger by calculating the distances from the contact anchors on the hand to different parts of the object. We then organize the contact information into language descriptors (e.g. "forefinger touches pump head and other fingers touch the bottle body."). Finally, we input both the brief guidance and the detailed contact information into the GPT3.5 to produce natural annotated guidance (e.g. "To use a lotion pump, press down on the pump head with your forefinger while holding the bottle with your other fingers.").
@@ 4 DexGYSGrasp framework
Progressive Grasp Objectives To address these challenges, we propose to decompose the complex learning objective into two more manageable objectives. The first objective is generative: it focuses on learning the grasp distribution, which does not prioritize quality but focuses on learning the grasp distribution with intention alignment and generation diversity. The second objective is regressive: it aims to refine the coarse grasp to a specific high-quality grasp with same intention. By decomposing the complex objectives, we reduce the learning difficulty of the generative objective as it does not concentrate on quality and avoids using penetration loss, which could interfere the learning process. Additionally, the learning of regression is less complex than distributions, as it merely requires adjusting the pose to a specific target within a small space. Hence, we can employ penetration loss to ensure that the refined dexterous hand avoids penetrating the object and with high quality.
@@ 4.2 Progressive Grasp Components
Quality Grasp Component. The generated grasps of the first component possess well-aligned intentions and high diversity, but suffer from poor grasp quality due to significant object penetration. Therefore, we introduce Quality Grasp Component to refine the grasp quality while maintaining intention consistency in a regressive manner. Specially, it takes the coarse pose $\hat{\mathcal{G}}^{dex}$, coarse hand point clouds $\mathcal{H}(\hat{\mathcal{G}}^{dex})$ and object point clouds $\mathcal{O}$ as input, and outputs the pose $\Delta\mathcal{G}^{dex}$. The refined grasp is obtained by $\tilde{\mathcal{G}}^{dex}=\hat{\mathcal{G}}^{dex}+\Delta\mathcal{G}^{dex}$. The training pairs of this component are constructed by collecting coarse grasps generated by the first component alongside the most similar ground-truth grasps that share the similar intentions. This ensures the training targets are aligned with the language intention, thereby guaranteeing that the refined grasps maintain consistency with the intended actions. Specially, QGC takes the coarse pose $\hat{\mathcal{G}}^{dex}$, coarse hand point clouds $\mathcal{H}(\hat{\mathcal{G}}^{dex})$ and object point clouds $\mathcal{O}$ as input, and outputs the pose $\Delta\mathcal{G}^{dex}$. The refined grasp is then given by $\tilde{\mathcal{G}}^{dex}=\hat{\mathcal{G}}^{dex}+\Delta\mathcal{G}^{dex}$.
@@ 5.1 Datasets and Evaluation Metrics
DexGYSNet Dataset / / 0.074 0.332 7.430 66.21 8.361
GraspCVAEcvae 3.138 0.096 0.054 0.551 0.179 1.762 0.179
GraspTTAgrasptta 12.19 0.111 0.071 0.188 2.111 6.150 3.869
SceneDiffuserscene_diffuser 1.679 0.045 0.083 0.253 0.346 3.455 0.387
DGTRxu2024dgtr 2.895 0.078 0.078 0.163 2.037 14.01 4.299
Ours 1.198 0.036 0.083 0.223 6.118 55.68 6.118
@@ 5.2 Implementation Details
For the construction of DexDYSNet, the step 2 and 3 are optimized for 20 and 300 iterations with learning rates of 0.01 and 0.0001 respectively. We set $\lambda^{1}_{pen}=100$ and set $\lambda^{1}_{spen}$, $\lambda^{1}_{joint}$, $\lambda^{1}_{cmap}$ each to 10. For training our framework, the training epochs are set to 100 for intention and diversity grasp component and 20 for Quality Grasp Component. The loss weights are configured as follows: $\lambda^{2}_{para}=\lambda^{3}_{para}=10$, $\lambda^{2}_{chamfer}=\lambda^{3}_{chamfer}=1$, $\lambda^{3}_{cmap}=10$, $\lambda^{3}_{pen}=100$, $\lambda^{3}_{spen}=10$. Throughout all training processes, the model is optimized with a batch size of 64 using the Adam optimizer, with a weight decay rate of $5.0\times 10^{-6}$. The initial learning rate is $2.0\times 10^{-4}$ and decay to $2.0\times 10^{-5}$ using a cosine learning rate sgdr scheduler. All experiment are implemented with PyTorch on a single RTX 4090 GPU.

### v2

@@ Abstract
This paper explores a novel task “Dexterous Grasp as You Say” (DexGYS), enabling robots to perform dexterous grasping based on human commands expressed in natural language. However, the development of this field is hindered by the lack of datasets with natural human guidance; thus, we propose a language-guided dexterous grasp dataset, named DexGYSNet, offering high-quality dexterous grasp annotations along with flexible and fine-grained human language guidance. Our dataset construction is cost-efficient, with the carefully-design hand-object interaction retargeting strategy, and the LLM-assisted language guidance annotation system. Equipped with this dataset, we introduce the DexGYSGrasp framework for generating dexterous grasps based on human language instructions, with the capability of producing grasps that are intent-aligned, high quality and diversity. To achieve this capability, our framework decomposes the complex learning process into two manageable progressive objectives and introduce two components to realize them. The first component learns the grasp distribution focusing on intention alignment and generation diversity. And the second component refines the grasp quality while maintaining intention consistency. Extensive experiments are conducted on DexGYSNet and real world environments for validation.
@@ 1 Introduction
In this paper, we explore a novel task, “Dexterous Grasp as You Say” (DexGYS), as shown in Figure 1. We can see that natural human guidance is provided in this task, and can be utilized to drive dexterous grasping generation, thereby facilitating more user-friendly human-robot interactions. However, the new task also brings in new challenges. First, the high costs of annotating dexterous pose and the corresponding language guidance, present a barrier for developing and scaling dexterous datasets. Second, the demands of generating dexterous grasps that ensure intention alignment, high quality and diversity, present considerable challenges to the model learning.
To address the first challenge, we propose a large-scale language-guided dexterous grasping dataset DexGYSNet. DexGYSNet is constructed in a cost-effective manner by exploiting human grasp behavior and the extensive capabilities of Large Language Models (LLM). Specially, we introduce the Hand-Object Interaction Retargeting (HOIR) strategy to transfer easily-obtained human hand-object interactions to robotic dexterous hand, to maintain contact consistency and high-quality grasp posture. Subsequently, we develop the LLM-assisted Language Guidance Annotation system to produce flexible and fine-grained language guidance for dexterous grasp data with the support of LLM. DexGYSNet dataset comprises 50,000 pairs of high-quality dexterous grasps and their corresponding language guidance, on 1,800 common household objects.
@@
With the support of the dataset, we now turn our way to overcome the second challenge. We propose the DexGYSGrasp framework for dexterous grasp generation, which aligns with intentions, ensures high quality, and maintains diversity. At the beginning, we find the difficulty of mastering all objectives simultaneously results from the commonly used penetration loss xu2024dgtr which used to avoid hand-object penetration. As shown in Figure 2, penetration loss substantially hinders the learning of grasp distribution, causing intention misalignment and reduced diversity. Conversely, despite the high diversity and aligned intention, the removal of penetration loss leads to unacceptable object penetration, making the grasp infeasible. Based on this finding, we design our DexGYSGrasp framework in a progressive strategy, decomposing the complex learning task into two sequential objectives managed by progressive components. Initially, the first component learns a grasp distribution, which focuses on intention consistency and diversity, optimizing effectively without the constraints of penetration loss. Subsequently, the second component refines the initial coarse grasps to high-quality ones with the same intentions and diversity. Our framework allows each component to focus on specific and manageable optimization objective, enhancing the overall performance of the generated grasps.
@@ 2.1 Dexterous Grasp Generation
Dexterous hand endows robots with the capability to manipulate objects in a human-like manner. Previous methods have achieved impressive results in ensuring grasp stability by analytical approaches analytic_1 ; DFC ; q1 ; dexgraspnet ; graspd and deep learning methods lu2023ugg ; unigrasp ; xu2024dgtr ; unidexgrasp ; efficientgrasp ; generating_multi_finger . However, the full potential of dexterous hands for intentional and human-like grasping has not been completely exploited in these methods. Recently, some works have focused on functional dexterous grasping chen2023TaskDex ; zhu2023FunctionDex2 ; wei2023FunctionDex1 ; zhu2021toward , aiming to achieve human-like capabilities that extend beyond grasp stability alone, but are still lack of flexibility and generalization. In this work, we explore a novel task, Language-guided Dexterous Grasp Generation, which fully leverages the dexterity of robotic hands and enable robot to execute dexterous grasp based on human natural language.
@@ 2.2 Grasp Datasets
The development of large-scale datasets has significantly improve the advancement of data-driven grasp methods, including parallel grasp fang20201billion ; eppner2021acronym ; economic ; tang2023task2_2 ; cai2024real , human grasp hasson2019obman ; chao2021dexycb ; yang2022oakink ; jian2023affordpose ; li2024semgrasp ; wang2024s2p , and dexterous grasp approaches ddg ; gendexgrasp ; chen2023TaskDex ; dexgraspnet ; unidexgrasp . Despite these advancements, the high cost of data collection remains a significant challenge, particularly in the domain of dexterous hands. Previous datasets for dexterous grasping primarily rely on physical analysis approaches DFC ; miller2004graspit to mitigate this issue. However, these approaches often lack the specific semantic context or corresponding language guidance necessary for constructing our language-guided dexterous task. In this paper, we present DexGYSNet dataset, with a cost-effective construction, providing high-quality dexterous grasp annotation along with flexible and fine-grained human language guidance.
@@ 2.3 Language-guided Robot Grasp
Language-guided robot grasp is important in robotics. Previous works focusing on parallel grippers have made strides in achieving task-oriented grasping murali2021task2_1 ; tang2023task2_2 ; tang2023graspgpt , language-guided grasping xu2023joint ; jin2024reasoning and manipulation jang2022bc ; mees2022matters ; driess2023palme ; shridhar2021cliport . In contrast to parallel grippers, dexterous hand boast a higher number of DOF (e.g., 28 for the Shadow Hand shadowhand ), enabling a broader dexterity. However, this high freedom also presents challenges for model learning. In this paper, we propose the DexGYSGrasp framework, capable of generating intention-aligned dexterous grasps with high-quality and diversity.
@@ 3.2 Hand-Object Interaction Retargeting
Our Hand-Object Interaction Retargeting (HOIR) aims to transfer human hand-object interaction to dexterous hand-object interaction as shown in Figure 3 . The source MANO romero2017MANO hand parameters are denoted as $\mathcal{G}^{m}\in{R}^{61}$. And the target dexterous hand parameters are denoted as $\mathcal{G}^{dex}$ = $(r,t,q)$, where $r\in\mathbf{SO(3)}$ represents the global rotation, $t\in\mathbb{R}^{3}$ is the translation in world coordinates, and $q\in\mathbb{R}^{J}$ is the joint angles for a $J$-DoF dexterous hand, for example $J=22$ for Shadow Handshadowhand .
Three steps are within the HOIR: pose initialization, fingertip alignment, and interaction refinement. In the first step, the dexterous poses are initialized by copying parameters from similar structures of human poses to establish better initial values. In the second step, the dexterous poses are optimized in the parameter space to align the fingertip positions $p_{k}^{dex,ft}$ with those of the human $p_{k}^{mano,ft}$. This achieves retargeting consistency, and the optimization objective can be formulated as follows:
@@
To improve the physical interaction feasibility while maintaining the consistency, the dexterous hand poses are further optimized in the third step by hand-object interaction and physical constraints losses dexgraspnet . Two key points are designed to maintain the consistency: preserving the contact area of the optimized pose consistent with the output from the second step, and keeping the translation fixed during this step. The optimization objective can be formulated as follows:
@@
Here, the object penetration loss $\mathcal{L}_{pen}$ penalizes the depth of hand-object penetration. The self-penetration loss $\mathcal{L}_{spen}$ penalizes the self-penetration. The joint angle loss $\mathcal{L}_{joint}$ penalizes the out-of-limit joint angles. The contact map loss $\mathcal{L}_{cmap}$ ensures the contact map on the object remains consistent with the output from the second stage. The details of losses can be found in Appendix A.1.5.
@@ 3.3 LLM-assisted Language Guidance Annotation
To annotate flexible and fine-grained language guidance for dexterous hand-object pairs with low-cost, we design a coarse-to-fine automated language guidance annotation system with the assistance of the LLM, inspired by cui2024anyskill ; li2024semgrasp , as shown in Figure 3. Specially, we initially generate brief guidance based on the object category and the brief human intention (e.g., "using a lotion pump"), which are collected by the human dataset yang2022oakink . Subsequently, we compile the contact information for each finger by calculating the distances from the contact anchors on the hand to different parts of the object. We then organize the contact information into language descriptors (e.g. "forefinger touches pump head and other fingers touch the bottle body."). Finally, we input both the brief guidance and the detailed contact information into the GPT3.5 to produce natural annotated guidance (e.g. "To use a lotion pump, press down on the pump head with your forefinger while holding the bottle with your other fingers."). More details about DexGYSNet construction can be found in Appendix A.2.
@@ 4 DexGYSGrasp framework
Progressive Grasp Objectives. To address these challenges, we propose to decompose the complex learning objective into two more manageable objectives. The first objective is generative: it focuses on learning the grasp distribution, which does not prioritize quality but focuses on learning the grasp distribution with intention alignment and generation diversity. The second objective is regressive: it aims to refine the coarse grasp to a specific high-quality grasp with same intention. By decomposing the complex objectives, we reduce the learning difficulty of the generative objective as it does not concentrate on quality and avoids using penetration loss which could interfere the learning process. Additionally, the learning of regression is less complex than distributions, as it merely requires adjusting the pose to a specific target within a small space. Hence, we can employ penetration loss to ensure that the refined dexterous hand avoids penetrating the object and with high quality.
@@ 4.2 Progressive Grasp Components
Quality Grasp Component. The generated grasps of the first component possess well-aligned intentions and high diversity, but suffer from poor grasp quality due to significant object penetration. Therefore, we introduce Quality Grasp Component to refine the grasp quality while maintaining intention consistency in a regressive manner. Specially, it takes the coarse pose $\hat{\mathcal{G}}^{dex}$, coarse hand point clouds $\mathcal{H}(\hat{\mathcal{G}}^{dex})$ and object point clouds $\mathcal{O}$ as input, and outputs the pose $\Delta\mathcal{G}^{dex}$. The refined grasp is obtained by $\tilde{\mathcal{G}}^{dex}=\hat{\mathcal{G}}^{dex}+\Delta\mathcal{G}^{dex}$. The training pairs of this component are constructed by collecting coarse grasps generated by the first component alongside the most similar ground-truth grasps that share the similar intentions. This ensures the training targets are aligned with the language intention, thereby guaranteeing that the refined grasps maintain consistency with the intended actions.
@@ 4.3 Progressive Grasp Loss
More details about loss function and model structure can be found in Appendix A.1.
@@ 5.1 Datasets and Evaluation Metrics
We split the DexDYSNet dataset at the object instance level, using 80% of the objects within each category for training and 20% for evaluation. Notably, none of the objects in the test set appear in the training set, ensuring that all experimental results are evaluated on unseen objects.
@@
$FID\downarrow$ $P\text{-}FID\downarrow$ $CD\downarrow$ $Con.\downarrow$ $Success\uparrow$ $Q_{1}\uparrow$ $Pen.\downarrow$ $\delta_{t}\uparrow$ $\delta_{r}\uparrow$ $\delta_{q}\uparrow$
GraspCVAEcvae 31.26 29.02 3.138 0.096 29.12% 0.054 0.551 0.179 1.762 0.179
GraspTTAgrasptta 35.41 33.15 12.19 0.111 43.46% 0.071 0.188 2.111 6.150 3.869
SceneDiffuserscene_diffuser 20.44 7.932 1.679 0.045 62.24% 0.083 0.253 0.346 3.455 0.387
DGTRxu2024dgtr 23.31 15.77 2.895 0.078 51.91% 0.078 0.163 2.037 14.01 4.299
Ours 6.538 5.595 1.198 0.036 63.31% 0.083 0.223 6.118 55.68 6.118
@@ 5.2 Implementation Details
For the construction of DexGYSNet, the step 2 and 3 are optimized for 20 and 300 iterations with learning rates of 0.01 and 0.0001 respectively. We set $\lambda^{1}_{pen}=100$ and set $\lambda^{1}_{spen}$, $\lambda^{1}_{joint}$, $\lambda^{1}_{cmap}$ each to 10. For training our framework, the training epochs are set to 100 for intention and diversity grasp component and 20 for Quality Grasp Component. The loss weights are configured as follows: $\lambda^{2}_{para}=\lambda^{3}_{para}=10$, $\lambda^{2}_{chamfer}=\lambda^{3}_{chamfer}=1$, $\lambda^{3}_{cmap}=10$, $\lambda^{3}_{pen}=100$, $\lambda^{3}_{spen}=10$. Throughout all training processes, the model is optimized with a batch size of 64 using the Adam optimizer, with a weight decay rate of $5.0\times 10^{-6}$. The initial learning rate is $2.0\times 10^{-4}$ and decay to $2.0\times 10^{-5}$ using a cosine learning rate sgdr scheduler. All experiment are implemented with PyTorch on a single RTX 4090 GPU.